# Topological phases of unitary dynamics: Classification in Clifford category

### v2

@@ 4. Time-reversal symmetry over qubits
Recall that the the multiplicative group of all finitely supported Pauli operators on ${\mathbb{Z}}^{\mathsf{d}}$, each of which is a tensor product of “clock shift” operator $\sum_{j\in{\mathbb{F}}_{p}}\ket{j+1}\bra{j}\in\operatorname*{{\mathsf{Mat}}}(p;{\mathbb{C}})$ and its Fourier transform $\sum_{j}e^{2\pi ij/p}\ket{j}\bra{j}\in\operatorname*{{\mathsf{Mat}}}(p;{\mathbb{C}})$, abelianizes to the additive group of Laurent polynomial “vectors” (column matrices) in ${\mathsf{d}}$ variables with coefficients in ${\mathbb{F}}_{p}$. Our results do not depend on particular generating set of the latter; the only exception so far was when we associated specific local Clifford ${\mathbb{C}}$-unitaries to elementary $\lambda$-unitaries, where the “clock shift” operators corresponded upper half blocks and their Fourier transforms lower half blocks.
@@
To prove the second statement, let $[V]\in{\mathfrak{U}}(R)$. Since ${\mathsf{U}}({\mathbb{F}}_{2})={\mathsf{EU}}({\mathbb{F}}_{2})$ and $({\mathfrak{u}}V)^{\dagger}\lambda({\mathfrak{u}}V)=\lambda$, we see that $[V]=[({\mathfrak{u}}V)^{-1}V]$ and ${\mathfrak{u}}(({\mathfrak{u}}V)^{-1}V)=I$. Thus, may assume that
@@ 5.6. Kernel of the descent maps from unitary groups
$\mu\zeta^{\dagger}$ =$=\begin{Bmatrix}L\\ L^{*}\end{Bmatrix}\begin{pmatrix}0&\pm I\\ \tfrac{1}{z}&0\end{pmatrix}\begin{Bmatrix}L^{*}\\ L\end{Bmatrix}:P^{*}\to P$
@@
(i) Suppose $s\in{\mathbf{z}}^{~{}\geq n}_{UB_{0}}\subset Q$. Obviously, ${\mathbf{z}}^{\geq 0}s=s$. Since $\eta\mp\eta^{\dagger}$ has $z$-degree zero and preserves ${\mathbf{z}}^{~{}\geq n}_{UB_{0}}$, we have $(\eta\mp\eta^{\dagger})s\in{\mathbf{z}}^{~{}\geq n}_{UB_{0}}$, Then, $\alpha^{\dagger}$ maps $(\eta\mp\eta^{\dagger})s$ into ${\mathbf{z}}^{>0}_{P_{0}}$ as seen by (71). Since both $\mu\zeta^{\dagger}$ and $\zeta^{\dagger}$ may decrease $z$-degree by at most one, we conclude that
@@
(iii) Suppose $s\in{\mathbf{z}}^{<0}_{B_{0}}\subset Q$. The map $\eta\mp\eta^{\dagger}$ preserves ${\mathbf{z}}^{<0}_{B_{0}}$ that is taken by $\alpha^{\dagger}$ to ${\mathbf{z}}^{<0}_{P_{0}}\subset P$. Then, both $\mu\zeta^{\dagger}$ and $\zeta^{\dagger}$ can only decrease $z$-degree. Therefore,
@@ 5.7. Going down from forms to unitaries
where in the top left entry, we used ${\mathbf{z}}^{<n}\Delta^{-1}{\mathbf{z}}^{<0}=\Delta^{-1}{\mathbf{z}}^{<0}$. Acting on $M\oplus M^{*}$, the projection $\nu$ does nothing. So, the diagonal terms act by identities, and the offdiagonal term is zero. Therefore, $F$ is an isomoprhism. ∎
@@
(107) $L(E^{\dagger}\Delta E,B,n+2n^{\prime})$ =$=\left\{\begin{pmatrix}E^{-1}a\\ {\mathbf{z}}^{\geq 0}E^{\dagger}\Delta a\end{pmatrix}+z^{n^{\prime}}\begin{pmatrix}E^{-1}b\\ E^{\dagger}\Delta b\end{pmatrix}+z^{n+n^{\prime}}\begin{pmatrix}{\mathbf{z}}^{<n^{\prime}}E^{-1}\Delta^{-1}c^{*}\\ E^{\dagger}c^{*}\end{pmatrix}\right\}$
@@
(110) ${\mathcal{X}}(E)\hat{L}$ =$=\left\{\begin{pmatrix}a\\ 0\end{pmatrix}\right\}\oplus z^{n^{\prime}}L(\Delta,B,n)\oplus\left\{\begin{pmatrix}0\\ c^{*}\end{pmatrix}\right\}\subseteq\bigoplus_{j=0}^{n+2n^{\prime}-1}z^{j}\begin{pmatrix}EB\\ E^{-\dagger}B^{*}\end{pmatrix}$
@@ 5.8. Going up from unitaries to forms
Although $\frac{1}{z-1}\notin{R[z,\tfrac{1}{z}]}$, by construction we know $z-1\in{R[z,\tfrac{1}{z}]}$ is a nonzerodivisor, and the formula makes sense in the one-element localization ${R[z,\tfrac{1}{z}]}_{z-1}$ and the final result is over ${R[z,\tfrac{1}{z}]}$. Since $U$ has a unit determinant, we see that ${\mathsf{B}}^{\uparrow}(U)$ has a unit determinant as well. It is easy to check that ${\mathsf{B}}^{\uparrow}(U)^{\dagger}=\pm{\mathsf{B}}^{\uparrow}(U)$.
@@
(129) ${\partial}_{n=1}(\Delta^{-1}B^{*},B)$ $=\begin{pmatrix}Q\\ 0\end{pmatrix}\oplus\begin{pmatrix}(1-z)I&0\\ 0&I\end{pmatrix}U^{-1}\begin{pmatrix}Q\\ 0\end{pmatrix}$
@@ 5.9. Kernel of the descent maps from Witt groups
We have to check if the image of $f$ is orthogonal to $S$. We look at the images of $a^{\prime\prime}$ and $b^{\prime\prime}$. First, $\Delta(z\Delta^{-1}c,a^{\prime\prime})-\Delta(\Delta^{-1}c,\Delta^{-1}(1-\nu+\tfrac{1}{z}\nu)\Delta a^{\prime\prime})=\Delta(\Delta^{-1}c,\tfrac{1}{z}a^{\prime\prime})-\Delta(\Delta^{-1}c,\Delta^{-1}(1-\nu+\tfrac{1}{z}\nu)\Delta a^{\prime\prime})=\Delta(\Delta^{-1}c,(\tfrac{1}{z}-1)\Delta^{-1}(1-\nu)\Delta a^{\prime\prime})=\mp(\tfrac{1}{z}-1)c(\Delta^{-1}(1-\nu)\Delta a^{\prime\prime})$, but $c\in{\mathbf{y}}^{\geq n}$ while $\Delta^{-1}(1-\nu)\circ\in projy^{<n}$. Hence, this expression has no $y^{0}$-term. Second, $\Delta^{-1}b^{\prime\prime}\in{\mathbf{y}}^{<n}$. Hence, $-\Delta(\Delta^{-1}c,\Delta^{-1}b^{\prime\prime})=\pm c(\Delta^{-1}b^{\prime\prime})$ has no $y^{0}$-term.
@@ 6. Coarse graining and blending equivalence
The surjectivity of ${\mathsf{B}}^{\downarrow}_{2\iota}$ is proved in 5.16. The injectivity of $\bar{\epsilon}$ is proved as in 5.18: if an ${R[z,\tfrac{1}{z}]}$-isomorphism $E$ shows an equivalence of $\bar{\epsilon}(z)(\Delta)$ to a trivial form for $[\Delta]\in{\mathfrak{W}}^{\mp}(R)$, then $\epsilon(E)$ (replacing $z$ by $1$) is an $R$-isomorphism showing that $\Delta$ is trivial. The composition being zero follows from that of 5.18 since ${\mathsf{S}}:{\widetilde{\mathfrak{W}}}^{\mp}{\mathfrak{W}}^{\mp}$ is surjective. To prove the exactness at the middle, let $[\Delta]\in{\mathfrak{W}}^{\mp}({R[z,\tfrac{1}{z}]})$ and suppose ${\mathsf{B}}^{\downarrow}_{2\iota}[\Delta]=0$. Borrowing the map $\hat{\mathsf{S}}$ in the proof of 5.18, we have $({\mathsf{B}}^{\downarrow}_{2\iota}\circ{\mathsf{S}})[\hat{\mathsf{S}}\Delta]=0$, implying that $[\hat{\mathsf{S}}\Delta]\in\bar{\epsilon}\,{\widetilde{\mathfrak{W}}}^{\mp}(R)$. Therefore, $[\Delta]={\mathsf{S}}[\hat{\mathsf{S}}\Delta]\in{\mathsf{S}}\bar{\epsilon}\,{\widetilde{\mathfrak{W}}}^{\mp}(R)=\bar{\epsilon}\,{\mathsf{S}}{\widetilde{\mathfrak{W}}}^{\mp}(R)=\bar{\epsilon}\,{\mathfrak{W}}^{\mp}(R)$. The sequence is split by $\epsilon$ and ${\mathsf{B}}^{\uparrow}_{2\iota-1}$. ∎
@@
We show that anything that is embedded into ${\mathfrak{L}}_{n}({\mathsf{d}},p)$ by some $\bar{\epsilon}$ can be ignored if we think of $\eta$- and $\lambda$-unitaries as translation invariant objects sitting on ${\mathbb{Z}}^{\mathsf{d}}$ modulo elementary unitaries that obey coarser translation invariance and think of quadratic and hermitian forms as translation invariant objects sitting on ${\mathbb{Z}}^{\mathsf{d}}$ modulo stable congruences that obey coarser translation invariance.

### v3

@@ 4. Time-reversal symmetry over qubits
Recall that the multiplicative group of all finitely supported Pauli operators on ${\mathbb{Z}}^{\mathsf{d}}$, each of which is a tensor product of “clock shift” operator $\sum_{j\in{\mathbb{F}}_{p}}\ket{j+1}\bra{j}\in\operatorname*{{\mathsf{Mat}}}(p;{\mathbb{C}})$ and its Fourier transform $\sum_{j}e^{2\pi ij/p}\ket{j}\bra{j}\in\operatorname*{{\mathsf{Mat}}}(p;{\mathbb{C}})$, abelianizes to the additive group of Laurent polynomial “vectors” (column matrices) in ${\mathsf{d}}$ variables with coefficients in ${\mathbb{F}}_{p}$. Our results do not depend on particular generating set of the latter; the only exception so far was when we associated specific local Clifford ${\mathbb{C}}$-unitaries to elementary $\lambda$-unitaries, where the “clock shift” operators corresponded upper half blocks and their Fourier transforms lower half blocks.
@@
To prove the second statement, let $[V]\in{\mathfrak{U}}(R)$. Since ${\mathsf{U}}({\mathbb{F}}_{2})={\mathsf{EU}}({\mathbb{F}}_{2})$ and $({\mathfrak{u}}V)^{\dagger}\lambda({\mathfrak{u}}V)=\lambda$, we see that $[V]=[({\mathfrak{u}}V)^{-1}V]$ and ${\mathfrak{u}}(({\mathfrak{u}}V)^{-1}V)=I$. Thus, we may assume that
@@ 5.6. Kernel of the descent maps from unitary groups
$\mu\zeta^{\dagger}$ =$=\begin{Bmatrix}L\\ L^{*}\end{Bmatrix}\begin{pmatrix}0&\pm I\\ \tfrac{1}{z}&0\end{pmatrix}\begin{Bmatrix}L^{*}\\ L\end{Bmatrix}:P^{*}\to P\,.$
@@
(i) Suppose $s\in{\mathbf{z}}^{~{}\geq n}_{UB_{0}}\subset Q$. Obviously, ${\mathbf{z}}^{\geq 0}s=s$. Since $\eta\mp\eta^{\dagger}$ has $z$-degree zero and preserves ${\mathbf{z}}^{~{}\geq n}_{UB_{0}}$, we have $(\eta\mp\eta^{\dagger})s\in{\mathbf{z}}^{~{}\geq n}_{UB_{0}}$, Then, $\alpha^{\dagger}$ maps $(\eta\mp\eta^{\dagger})s$ into ${\mathbf{z}}^{>0}_{P_{0}}$ as seen by (71). Since both $\mu\zeta^{\dagger}$ and $\zeta^{\dagger}$ may decrease the $z$-degree by at most one, we conclude that
@@
(iii) Suppose $s\in{\mathbf{z}}^{<0}_{B_{0}}\subset Q$. The map $\eta\mp\eta^{\dagger}$ preserves ${\mathbf{z}}^{<0}_{B_{0}}$ that is taken by $\alpha^{\dagger}$ to ${\mathbf{z}}^{<0}_{P_{0}}\subset P$. Then, both $\mu\zeta^{\dagger}$ and $\zeta^{\dagger}$ can only decrease the $z$-degree. Therefore,
@@ 5.7. Going down from forms to unitaries
where in the top left entry, we used ${\mathbf{z}}^{<n}\Delta^{-1}{\mathbf{z}}^{<0}=\Delta^{-1}{\mathbf{z}}^{<0}$. Acting on $M\oplus M^{*}$, the projection $\nu$ does nothing. So, the diagonal terms act by identities, and the offdiagonal term is zero. Therefore, $F$ is an isomorphism. ∎
@@
(107) $L(E^{\dagger}\Delta E,B,n+2n^{\prime})$ =$=\left\{\begin{pmatrix}E^{-1}a\\ {\mathbf{z}}^{\geq 0}E^{\dagger}\Delta a\end{pmatrix}+z^{n^{\prime}}\begin{pmatrix}E^{-1}b\\ E^{\dagger}\Delta b\end{pmatrix}+z^{n+n^{\prime}}\begin{pmatrix}{\mathbf{z}}^{<n^{\prime}}E^{-1}\Delta^{-1}c^{*}\\ E^{\dagger}c^{*}\end{pmatrix}\right\}\,.$
@@
(110) ${\mathcal{X}}(E)\hat{L}$ =$=\left\{\begin{pmatrix}a\\ 0\end{pmatrix}\right\}\oplus z^{n^{\prime}}L(\Delta,B,n)\oplus\left\{\begin{pmatrix}0\\ c^{*}\end{pmatrix}\right\}\subseteq\bigoplus_{j=0}^{n+2n^{\prime}-1}z^{j}\begin{pmatrix}EB\\ E^{-\dagger}B^{*}\end{pmatrix}\,.$
@@ 5.8. Going up from unitaries to forms
Although $\frac{1}{z-1}\notin{R[z,\tfrac{1}{z}]}$, by construction we know $z-1\in{R[z,\tfrac{1}{z}]}$ is a non-zero-divisor, and the formula makes sense in the one-element localization ${R[z,\tfrac{1}{z}]}_{z-1}$ and the final result is over ${R[z,\tfrac{1}{z}]}$. Since $U$ has a unit determinant, we see that ${\mathsf{B}}^{\uparrow}(U)$ has a unit determinant as well. It is easy to check that ${\mathsf{B}}^{\uparrow}(U)^{\dagger}=\pm{\mathsf{B}}^{\uparrow}(U)$.
@@
(129) ${\partial}_{n=1}(\Delta^{-1}B^{*},B)$ $=\begin{pmatrix}Q\\ 0\end{pmatrix}\oplus\begin{pmatrix}(1-z)I&0\\ 0&I\end{pmatrix}U^{-1}\begin{pmatrix}Q\\ 0\end{pmatrix}\,.$
@@ 5.9. Kernel of the descent maps from Witt groups
We have to check if the image of $f$ is orthogonal to $S$. We look at the images of $a^{\prime\prime}$ and $b^{\prime\prime}$. First, $\Delta(z\Delta^{-1}c,a^{\prime\prime})-\Delta(\Delta^{-1}c,\Delta^{-1}(1-\nu+\tfrac{1}{z}\nu)\Delta a^{\prime\prime})=\Delta(\Delta^{-1}c,\tfrac{1}{z}a^{\prime\prime})-\Delta(\Delta^{-1}c,\Delta^{-1}(1-\nu+\tfrac{1}{z}\nu)\Delta a^{\prime\prime})=\Delta(\Delta^{-1}c,(\tfrac{1}{z}-1)\Delta^{-1}(1-\nu)\Delta a^{\prime\prime})=\mp(\tfrac{1}{z}-1)c(\Delta^{-1}(1-\nu)\Delta a^{\prime\prime})$, but $c\in{\mathbf{y}}^{\geq n}$ while $\Delta^{-1}(1-\nu)\circ\in{\mathbf{y}}^{<n}$. Hence, this expression has no $y^{0}$-term. Second, $\Delta^{-1}b^{\prime\prime}\in{\mathbf{y}}^{<n}$. Hence, $-\Delta(\Delta^{-1}c,\Delta^{-1}b^{\prime\prime})=\pm c(\Delta^{-1}b^{\prime\prime})$ has no $y^{0}$-term.
@@ 6. Coarse graining and blending equivalence
The surjectivity of ${\mathsf{B}}^{\downarrow}_{2\iota}$ is proved in 5.16. The injectivity of $\bar{\epsilon}$ is proved as in 5.18: if an ${R[z,\tfrac{1}{z}]}$-isomorphism $E$ shows an equivalence of $\bar{\epsilon}(z)(\Delta)$ to a trivial form for $[\Delta]\in{\mathfrak{W}}^{\mp}(R)$, then $\epsilon(E)$ (replacing $z$ by $1$) is an $R$-isomorphism showing that $\Delta$ is trivial. The composition being zero follows from that of 5.18 since ${\mathsf{S}}:{\widetilde{\mathfrak{W}}}^{\mp}\to{\mathfrak{W}}^{\mp}$ is surjective. To prove the exactness at the middle, let $[\Delta]\in{\mathfrak{W}}^{\mp}({R[z,\tfrac{1}{z}]})$ and suppose ${\mathsf{B}}^{\downarrow}_{2\iota}[\Delta]=0$. Borrowing the map $\hat{\mathsf{S}}$ in the proof of 5.18, we have $({\mathsf{B}}^{\downarrow}_{2\iota}\circ{\mathsf{S}})[\hat{\mathsf{S}}\Delta]=0$, implying that $[\hat{\mathsf{S}}\Delta]\in\bar{\epsilon}\,{\widetilde{\mathfrak{W}}}^{\mp}(R)$. Therefore, $[\Delta]={\mathsf{S}}[\hat{\mathsf{S}}\Delta]\in{\mathsf{S}}\bar{\epsilon}\,{\widetilde{\mathfrak{W}}}^{\mp}(R)=\bar{\epsilon}\,{\mathsf{S}}{\widetilde{\mathfrak{W}}}^{\mp}(R)=\bar{\epsilon}\,{\mathfrak{W}}^{\mp}(R)$. The sequence is split by $\epsilon$ and ${\mathsf{B}}^{\uparrow}_{2\iota-1}$. ∎
@@
We are going to show that anything that is embedded into ${\mathfrak{L}}_{n}({\mathsf{d}},p)$ by some $\bar{\epsilon}$ can be ignored if we consider coarser translation invariance.
@@
Declarations
Data availability. Data availability does not apply to this research since no data was used or generated.
Competing Interests. The author declares that there is no conflict of interest.